OpenStudy (waheguru):

How would you factor this 0= x^2+x-2

OpenStudy (mertsj):

\[(x+2)(x-1)\]

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\]\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]\[(x-x_1)(x-x_2)=0\] _____ \[x^2+x-2=0\]\[x_{1,2}=\frac{-1\pm\sqrt{1^2-4\times 1\times -2}}{2\times1}\\ \qquad=\frac{-1\pm\sqrt{1+8}}{2}\\ \qquad=\frac{-1\pm\sqrt 9}{2}\\ \qquad=\frac{-1\pm3}{2}\\ \qquad=-\frac42,\frac{2}2\\ \qquad=-2,1\] \[(x+2)(x-1)=0\]